OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will always result in the same number. You will use the inverse function that you created in problem number 5 for g(x).

OpenStudy (anonymous):

My inverse function - x-6 ------ = f(x) 10 I was learning it off of - http://openstudy.com/study#/updates/52331ea2e4b0af32a07890c2

OpenStudy (anonymous):

@E.ali This is what I needed help on!

OpenStudy (anonymous):

@Yttrium Would ya help me out?

ganeshie8 (ganeshie8):

you're not @Brad1996 eh ?

OpenStudy (anonymous):

@ganeshie8 No, i was learning off of his question though

ganeshie8 (ganeshie8):

okie nice :)

OpenStudy (anonymous):

@ganeshie8 Thanks, I'm just really confused, I would like if you solved it step by step and showed me in details how you did it.. I don't understand this at all and I'm falling behind due to it

ganeshie8 (ganeshie8):

okay lets do it... read the question quickly, and can u write down the f(x) and g(x) ?

OpenStudy (anonymous):

@ganeshie8 f(x) = 10x+ 6 g(x) = (x-6)/10 if I remember correctly.

ganeshie8 (ganeshie8):

Awesome ! yes they're the two functions, which are inverses of eachother

ganeshie8 (ganeshie8):

next, Assign any number to x.

ganeshie8 (ganeshie8):

wats ur fac number ? :)

ganeshie8 (ganeshie8):

*fav number

OpenStudy (anonymous):

I'll choose the number 2. lol

ganeshie8 (ganeshie8):

very smart :) cuz 2 is small ha

ganeshie8 (ganeshie8):

ok, x = 2

ganeshie8 (ganeshie8):

next, Using complete sentences, explain whether f(g(x)) and g(f(x)) will always result in the same number.

ganeshie8 (ganeshie8):

to do this, we need to calculate two things :- f(g(2)) and g(f(2))

OpenStudy (anonymous):

Hmm alright, and how would we do that?

ganeshie8 (ganeshie8):

lets do f(g(2)) first f(g(2)) ^

ganeshie8 (ganeshie8):

do u see g(2) is inside, so we need to calculate it first

ganeshie8 (ganeshie8):

g(2) = ?

OpenStudy (anonymous):

Well, I'm assuming we would multiply them which would be  since there doesn't seem to be much else you can do.

ganeshie8 (ganeshie8):

i see wat u did there, but here \(g\) is not a variabel, \(g\) is a function !!!! \(\large g(x) = \frac{x-6}{10}\)

ganeshie8 (ganeshie8):

so, \(\large g(2) = \frac{2-6}{10}\)

ganeshie8 (ganeshie8):

simplify,

ganeshie8 (ganeshie8):

above, basically, to get g(2), we just replaced x wid 2

OpenStudy (anonymous):

Simplified would be g(2) = -2/5 

OpenStudy (anonymous):

Really bad computer drawing ability, lol.

ganeshie8 (ganeshie8):

that looks good, sorry, i gtg but il be back in 20 mibutes

OpenStudy (anonymous):

Okay, I'll work on another class while I'm waiting. :)

ganeshie8 (ganeshie8):

m back @Smores

ganeshie8 (ganeshie8):

u still here ? :)

OpenStudy (anonymous):

Yes I'm back now @ganeshie8

ganeshie8 (ganeshie8):

so where are we.. :)

ganeshie8 (ganeshie8):

we're trying to find :- f(g(2)) and g(f(2))

ganeshie8 (ganeshie8):

for f(g(2)) you have already found g(2) = -2/5 right ?

OpenStudy (anonymous):

Can somebody please give me the answer? I am falling behind in Math! D:

ganeshie8 (ganeshie8):

i can help u if u want :)

OpenStudy (anonymous):

@ganeshie8 That's fine but I can't wait much longer, I really need the answer...

ganeshie8 (ganeshie8):

uve disappeared halfway in between 4 days ago, and al of a sudden u want answer and cant wait -_-

ganeshie8 (ganeshie8):

i almost forgot about this problem ! if u could tell me where u got stuck i can give a hand... i dont mean to be offensice, just saying : dont expect help from me without you trying it first

OpenStudy (anonymous):

@ganeshie8 I have tried, I can't do it.

ganeshie8 (ganeshie8):

still stuck on this ? :)

OpenStudy (anonymous):

@ganeshie8 Yes and I've already tried to no avail.

ganeshie8 (ganeshie8):

its okay, lets work it and finish

ganeshie8 (ganeshie8):

if i remember correctly, we're trying to do below :- ``` to do this, we need to calculate two things :- f(g(2)) and g(f(2)) ```

ganeshie8 (ganeshie8):

wat didi u get for f(g(2)) ?

ganeshie8 (ganeshie8):

here are teh functions for reference :- ``` f(x) = 10x+ 6 g(x) = (x-6)/10 ```

OpenStudy (anonymous):

@ganeshie8 I'm honestly not sure how I'd solve that.

ganeshie8 (ganeshie8):

its okay, once u see me doing it, u will understand

ganeshie8 (ganeshie8):

to find, f(g(2)) first find g(2) g(x) = (x-6)/10 g(2) = ?

ganeshie8 (ganeshie8):

simply replace x wid 2, and simplify

ganeshie8 (ganeshie8):

g(x) = (x-6)/10 g(2) = (2-6)/10 = ?

ganeshie8 (ganeshie8):

i want u simplify it